OpenStudy (anonymous):

factor... x^11(x+12)+10(x+12)

OpenStudy (anonymous):

x^(12) + 12x^(11) +10x+120 = (x^(11) + 10)(x + 12)

OpenStudy (anonymous):

what times what gives 120 the think when I add x values what will I get

OpenStudy (anonymous):

am i using 12 because its common in both factors?

OpenStudy (anonymous):

you are using 12 because 12 * 10 = 120

OpenStudy (anonymous):

also think about it 12*x^(11) = 12x^(11) and 10 * x = 10x

OpenStudy (anonymous):

i think that is the way i worked it, but i still came up with the same answer as you did. thanks for the help.

OpenStudy (anonymous):

meh factoring isn't hard you just need to develop the thought process

OpenStudy (anonymous):

yea. what i did was took the common factors and used them. I think my problem is trying to find the "short" ways to figure out problems. most of the time i get them correct lol